5x-2y=-8
3x-8y=-32 how do you do this

Answer by unlockmath(1688)

Hello,
First we want to get it so we can cancel either the x or y when we add both equations together. So let's multiply the first equation by -4 to give us:
-20x+8y=+32 Now we can add this to the second equation.
3x-8y=-32
We get -17x=0 so x = 0
Now plug 0 in for x in either original equation and we get:
5(0)-2y=-8
y=4
To check these put them into both equations and it should prove correct.
RJ Toftness
www.math-unlock.com